OpenStudy (anonymous):

Help Please! Find the sum of the following infinite geometric series, if it exists. 1/2+1/4+1/8+1/16+...

OpenStudy (anonymous):

1/32

OpenStudy (anonymous):

how do you get that

OpenStudy (anonymous):

If you divide each one by 2 the next sequence is next like 1/2 divided by 2 = 1/4 1/4 divided by 2 equals = 1/8 divided by 2 =1/16 and 1/16 divided by 2 = 1/32

OpenStudy (anonymous):

theyll be in decimals if you use a calculator thats all

OpenStudy (anonymous):

thats the sum?

OpenStudy (anonymous):

no its division but 1/32 is the next sequence

OpenStudy (anonymous):

and to find the next sequence after 1/32 just put 1/32 divided by 2 which is 0.015625 but in fraction its 1/64

OpenStudy (anonymous):

I need to know how to find the sum of them series.. Do you just add them?

OpenStudy (anonymous):

yeah you can add the last number like 2+2 = 4 and 4+4 = 8

OpenStudy (anonymous):

8+8 = 16 and 16 + 16 = 32

OpenStudy (anonymous):

that can also be a way you can do it

OpenStudy (anonymous):

nevermind you dont understand thanks anyway

OpenStudy (anonymous):

k you welcome